# Depression Detection on Social Media with Large Language Models

### v1

@@ 4.1.1 Implementation Details
In our implementation, following orabi2018deep ; cai2023depression , we utilize cosine similarity to calculate the similarity between embeddings. To enhance the usability of our method in low-compute resource settings, we employ the low-resource-demanding pre-trained embedding model, gte-small-zh  [^5] https://huggingface.co/thenlper/gte-small, which operates smoothly with just 1GB of memory. The embedding model in our system can easily be switched to other higher-performance models to further improve performance. For the LLM, we use GPT-3.5-Turbo-1103  [^6] https://platform.openai.com/docs/models/gpt-3-5-turbo, which requires only an internet connection to interact with it through the API service provided by OpenAI. The LLM in our system can also be substituted with open-source models, such as Mentallama yang2023mentalllama , to further reduce monetary costs. Overall, our system can run at low computational costs, enhancing its accessibility. We utilize XGBoost chen2016xgboost for an efficient implementation of Gradient Boosting Trees.
4.1.2 Dataset
We utilize a large-scale, open-source dataset, SWDD cai2023depression to suppor our experiments. After expert screening, we keep 1000 depressed users and 19000 normal users, to simulate the actual depression rate. The dataset statistics are shown in Table 2. Training, validation, and test sets were divided in a 7:1:2 ratio, as recommended by wu2023exploring . We give a detailed discussion of the dataset in Section A.
@@ 4.1.2 Dataset
Table 2: Statistics of our utilized dataset.
Category Method Precision Recall F1-score AUROC AUPRC
Traditional Method TF-IDF+XGBoost 0.3644 0.4300 0.3945 0.9023 0.4303
Deep Learning-Based Methods HAN 0.5702 0.6500 0.6075 0.8929 0.5864
Mood2Content 0.7216 0.7000 0.7106 0.9537 0.7774
PLM-Based Methods FastText 0.7467 0.5600 0.6400 0.9441 0.6255
gte-small 0.6359 0.6526 0.6200 0.9499 0.6959
BERT 0.6667 0.6400 0.6531 0.9481 0.7102
MentalRoBERTa 0.7326 0.6300 0.6774 0.9423 0.6880
LLM-Based Methods ChatGPT 0.0875 0.7100 0.1559 0.6603 0.0767
MentalLLama 0.0899 0.7800 0.1612 0.6821 0.0811
Our Method DORIS 0.7596 0.7900 0.7596 0.9715 0.8134
Table 3: Performance of DORIS and baselines. The best scores are in bold, and second best scores are underlined.
4.1.3 Baseline Methods
We employed various baselines, including methods combining traditional feature extraction with classifiers: TF-IDF+XGBoost ramos2003using ; chen2016xgboost ; wu2023exploring , deep learning approaches: HAN yang2016hierarchical , PLM-based methods: FastText joulin2016fasttext , BERT devlin2018bert , MentalRoBERTa ji2023domain , and Mood2Content cai2023depression , as well as large language model methods like MentalLLama yang2023mentalllama and ChatGPT. For LLM-based approaches, we design prompts following yang2023towards . Due to the token window constraint of large models, in some instances, we truncated the data to retain as many recent tweets as possible while ensuring the total prompt length did not exceed the limit.
4.1.4 Evaluation Metrics
Following prior works, we evaluate our method using five metrics: Precision, Recall, F1, AUROC, and AUPRC. Precision measures the proportion of true positive predictions in all positive predictions. Recall assesses the proportion of true positive predictions out of all actual positives. F1 Score is the harmonic mean of Precision and Recall, providing a balance between them. AUROC measures the probability that a classifier will rank a random positive instance higher than a random negative one. AUPRC is the area under the curve plotting precision against recall under various thresholds. It reflects the classifier’s ability to identify positive instances among a set. For depression detection, AUPRC can be considered as the most important metric, as it best reflects the classifier’s performance on highly imbalanced datasets davis2006relationship .
4.2 Experimental Results
4.2.1 Overall Performance
In Table 3, we present a performance comparison between our method and comparison methods. The results highlight the superior performance of our approach. Specifically, we observe:
  • Our method outperforms all baselines across all metrics. On AUPRC, which best reflects model performance on imbalanced datasets, DORIS shows an absolute improvement of 0.0360 over the best-performing baseline, Mood2Content. Additionally, in our implementation, we utilized gte-small and ChatGPT as the backbone. Compared to these models, our method significantly improves classification performance, demonstrating the effectiveness of our design.
  • Models specifically designed for depression detection tasks outperform general models. Among various models, those designed for depression detection or mental health analysis, namely Mood2Content, MentalRoBERTa, and MentalLLama, achieved better performance. Our DORIS method, integrating medical understanding of depression and actual diagnostic practices, achieved the best results through task-specific design.
  • Direct classification with LLMs yields low performance. Our experiments show that both ChatGPT and MentalLLama underperformed, lagging significantly behind DORIS and other baselines on the crucial AUPRC metric. Methods based on LLMs exhibit high Recall but lower performance on other metrics, primarily due to their tendency to classify users as having depression. In our experiments, some truncation of input to LLMs was necessary, yet this truncation is not the cause of the poor performance of LLM-based methods, as it does not alter their inclination to predict the positive class. Our approach effectively harnesses the capabilities of LLMs without directly relying on their output for judgments, thereby achieving optimal results.
4.2.2 Ablation Study
In our method, the GBT classifier utilizes three constructed features to produce the final prediction results. Here we individually remove each of these feature sets from the classifier’s input and observe the corresponding performance changes. By this means, we aim to validate the contribution of these features to the overall performance gain. The experimental results are presented in Table 4.
F1-score AUROC AUPRC
Full Design 0.7596 0.9715 0.8134
w/o DC Feature 0.6867 0.9679 0.7739
w/o MC Representation 0.7415 0.9660 0.7932
w/o PH Representation 0.7200 0.9660 0.7817
Table 4: Experimental results of ablation study. DC denotes diagnostic criteria, MC denotes mood course, and PH denotes post history.
It is evident that removing any one of these three features results in a significant decrease in our system’s performance. The most notable decline occurs upon removing the diagnostic criteria feature, with AUPRC dropping to 0.7739, which is lower than Mood2Content’s 0.7774. This highlights the importance of the diagnostic criteria feature. Similarly, the performance significantly declines upon the removal of the other two features, proving the effectiveness of every part of DORIS’s design.
4.2.3 Hyperparameter Study
(a) The impact of $k$.
(b) The impact of $m$.

### v2

@@ 3.1 Implementation Details
Following Orabi et al. (2018), we utilize cosine similarity to calculate the similarity between embeddings. To enhance the usability of our method in low-compute resource settings, we employ a low-resource-demanding pre-trained text embedding model, gte-small Li et al. (2023), which operates smoothly with just 1GB of memory. The text embedding model in our system can easily be switched to other higher-performance models to further improve performance. For the LLM, we use GPT-4o mini  [^1] https://platform.openai.com/docs/models/GPT-4o-mini, which requires only an internet connection to interact with it through the API service provided by OpenAI. We have verified that OpenAI’s terms of service prohibit storing or retaining any user-provided text data, ensuring data privacy. The LLM in our system can also be substituted with open-source models deployed locally, such as Mentallama (Yang et al., 2023b), to further ensure data privacy. Overall, our system can run at low computational costs, enhancing its accessibility. We utilize XGBoost (Chen and Guestrin, 2016) for an efficient implementation of Gradient Boosting Trees. We run each experiment 5 times and report the average performance across these runs.
Table 2: Performance of DORIS and baselines on the SWDD dataset. The best scores are in bold, and second best scores are underlined.
Category Method Precision Recall F1-score AUROC AUPRC
Traditional Method TF-IDF+XGBoost 0.3644 0.4312 0.3945 0.9023 0.4303
Deep Learning-Based Methods HAN 0.5702 0.6524 0.6075 0.8929 0.5864
Mood2Content 0.7216 0.6996 0.7106 0.9537 0.7774
[…]
gte-small 0.7359 0.6524 0.6916 0.9499 0.6959
BERT 0.6667 0.6294 0.6531 0.9481 0.7102
MentalRoBERTa 0.7326 0.6272 0.6774 0.9423 0.6880
[…]
PHQ9 (Vector) 0.7221 0.6852 0.7032 0.9479 0.7631
LLM-Based Methods GPT-4o mini 0.0895 0.7122 0.1590 0.6603 0.0767
[…]
Our Method DORIS 0.7606 0.7902 0.7750 0.9722 0.8147
3.2 Dataset
Our primary study utilizes a cohort from the ethically-sourced SWDD dataset Cai et al. (2023), consisting of 1,000 expert-identified depressed users and 19,000 control users to reflect a real-world 1:19 prevalence ratio. Following best practices Eichstaedt et al. (2018) and accommodating model constraints, we retained only the final six months of each user’s post history. This resulted in a corpus of approximately 70k posts from the depressed group and 1.3M from the control group. To further validate the generalizability of our approach, we also conducted experiments on the Twitter Mental Disorder dataset  [^2] https://www.kaggle.com/datasets/rrmartin/twitter-mental-disorder-tweets-and-musics.. For a more detailed description of the datasets, please refer to Appendix B.
3.3 Baseline Methods
We employ various baselines, including methods combining traditional feature extraction with classifiers: TF-IDF+XGBoost (Ramos et al., 2003; Chen and Guestrin, 2016; Wu et al., 2023), deep learning approaches: HAN (Yang et al., 2016), Mood2Content (Cai et al., 2023), and AMM-Net (Sarkar et al., 2022), PLM-based methods: FastText (Joulin et al., 2016), BERT (Devlin et al., 2018), gte-small Li et al. (2023) and MentalRoBERTa (Ji et al., 2023), medical knowledge-guided methods: PHQ9 (Score) and PHQ9 (Vector) Nguyen et al. (2022), as well as LLM-based methods like MentalLLama (Yang et al., 2023b) and GPT-4o mini. A more detailed description of comparison methods is in Appendix C.
3.4 Evaluation Metrics
Following prior works, we evaluate our method using five metrics: Precision, Recall, F1, AUROC, and AUPRC. AUPRC can be considered as the most important metric, as it best reflects the classifier’s performance on highly imbalanced datasets (Davis and Goadrich, 2006).
4 Experimental Analysis
We conducted extensive experiments to address the following research questions:
  • RQ1: How does our proposed system, DORIS, perform on the depression detection task compared to state-of-the-art methods?
  • RQ2: How effective is DORIS at generating clinically relevant and evidence-based explanations for its predictions?
  • RQ3: To what extent does each component of DORIS contribute to its overall performance?
  • RQ4: How do key hyperparameter settings impact the performance of our method?
Due to space constraints, we present the experimental results for RQ4 in the Appendix.
4.1 Overall Performance (RQ1)
The results in Table 2 demonstrate the superiority of our approach on the SWDD dataset. DORIS outperforms all baselines across all metrics, achieving an absolute improvement of 0.0361 in AUPRC over the strongest baseline, AMM-Net. This performance gain stems from our hybrid design, which addresses the weaknesses of alternative methods.
Our experiments confirm that directly using LLMs like GPT-4o mini for classification leads to poor performance, exhibiting high recall but very low precision. Furthermore, while other methods incorporate medical knowledge (e.g., PHQ9-based models), their reliance on less capable backbones like BERT limits their ability to understand complex symptom descriptions. In contrast, DORIS leverages a powerful LLM to effectively model clinical criteria and mood courses. By using the LLM for sophisticated, clinically-grounded feature engineering rather than direct prediction, our task-specific design effectively combines the strengths of modern language models and robust classifiers, leading to state-of-the-art results.
We also conduct experiments on the Twitter Mental Disorder dataset to further validate the effectiveness and generalizability of DORIS. The results are presented in Table 3.
Table 3: Performance comparison on the Twitter Mental Disorder Dataset.
Model F1-score AUROC AUPRC
TF-IDF+XGBoost 0.9156 0.7149 0.9675
FastText 0.9189 0.7879 0.9770
Mood2Content 0.9101 0.8112 0.9775
gte-small 0.9028 0.8089 0.9812
BERT 0.9211 0.7581 0.9753
MentalRoBERTa 0.9211 0.8143 0.9824
DORIS 0.9296 0.8647 0.9878
As the results show, DORIS consistently outperforms strong baselines on this additional dataset across all metrics. This reinforces the findings from the SWDD dataset and demonstrates the robustness of our approach across different data sources.
4.2 DORIS’s explainability (RQ2)
In our system, we annotate symptoms exhibited in users’ posts, summarize the users’ mood course, and based on these, provide explanations for the system’s final decision output. An example is presented in Figure 3 to aid in better understanding the operation of our system. As seen, during the diagnostic criteria feature construction stage, our method accurately identifies depression-related symptoms in posts. In the mood course representation construction stage, our approach selects posts with high emotional density, thereby generating a mood course description. By combining identified symptoms of depression and the user’s mood course, the system can generate evidence-supported explanations for its judgment. The DORIS system’s accurate annotations of depression symptoms, detailed analysis of mood course, and the final evidence-supported explanations are all visible to mental health professionals. This allows the system to serve as a supportive clinical tool in depression assessment.
Figure 3: A case study of DORIS’s output.
4.3 Ablation Study (RQ3)
Table 4: Experimental results of ablation study.
F1-score AUROC AUPRC
Full Design 0.7750 0.9722 0.8147
w/o Diagnostic Criteria Feature 0.6867 0.9679 0.7739
w/o LLM Annotation 0.7028 0.9685 0.7861
w/o Mood Course Representation 0.7415 0.9660 0.7932
w/o Mood Course Summary 0.7635 0.9699 0.8035
w/o Post History Representation 0.7345 0.9660 0.7817
In our method, the GBT classifier utilizes three constructed features: diagnostic criteria feature, mood course representation, and post history representation. We conduct ablation studies by removing each feature and observing the performance changes. The results are presented in Table 4.
The results clearly indicate that the diagnostic criteria feature is the most critical component. Its removal causes the most substantial performance drop, with AUPRC decreasing from 0.8147 to 0.7739. This underscores the crucial role of integrating established medical knowledge. Within this module, the LLM-based symptom annotation is also vital; replacing it with a simpler template-matching approach (w/o LLM Annotation) still results in a significant AUPRC drop to 0.7861, demonstrating the LLM’s superior ability to capture nuanced symptom expressions.
The post history representation, which provides a holistic view of a user’s activity, is also highly impactful. Removing it leads to the second-largest performance decline (AUPRC of 0.7817), confirming its importance in capturing contextual information that more targeted features might miss. Finally, the mood course representation proves essential for modeling temporal emotional patterns. Its absence reduces the AUPRC to 0.7932, with its sub-component, the LLM-generated summary, also showing a clear contribution.
These results collectively demonstrate that each component of DORIS makes meaningful contributions to its overall performance, confirming the effectiveness of every part of DORIS’s design.
5 Conclusion and Future Work
In this work, we present DORIS, a novel depression detection system that achieves high accuracy and interpretability by aligning with clinical diagnostic practices. The system integrates the widely used DSM-5 diagnostic criteria with analyses of mood course to deliver robust and explainable results. Extensive experiments validate the effectiveness of our method and the contribution of each design component. Future work will focus on pathways for real-world application, including integration with telehealth platforms and pilot studies within clinical settings to assess deployment efficacy and scalability.
Limitations
This work’s primary limitation is its reliance on a focused set of datasets. While we validated our approach on two different datasets, further validation across more diverse data sources is necessary to ensure broader generalizability. Furthermore, while our experimental setup simulates a realistic pre-screening scenario (e.g., by using a 1:19 imbalance ratio), validation in a real-world clinical setting is the gold standard and a necessary next step for deployment. Finally, future research should extend this approach to other mental health conditions, such as anxiety and bipolar disorder, to broaden its clinical relevance and impact.
Ethics Statement
Our work on depression detection, a sensitive task, was guided by a strong ethical framework. We detail our primary considerations and mitigation strategies below.
Research Positioning and Intended Use.
We position this work primarily as a Text Mining study aimed at extracting signals indicative of depression from unstructured text, aligning with established clinical criteria (DSM-5). We emphasize that this system is designed strictly as a supportive tool for qualified mental health professionals. It is not an autonomous diagnostic system and should never be deployed directly to end-users or used to make final clinical decisions. The explainable evidence provided (e.g., symptom annotations) is intended to help professionals verify risk signals. This "human-in-the-loop" model is key to mitigating the risks of potential model bias and error inherent in LLMs. Any real-world implementation must be supervised by practitioners and preceded by thorough validation from clinical experts.
Data Privacy and Security.
All data was sourced from users who provided explicit consent for their content to be used in mental health research. To protect user privacy, we moderately obfuscated all examples in this paper per established guidelines (Bruckman, 2002). While our research utilized a commercial API, we have engineered the system to be deployable with secure, locally-hosted open-source LLMs, thereby eliminating the need to share sensitive patient data with third-party services in a clinical setting.
Risks of Misuse and Mitigation.
We recognize the potential for misuse, such as exploiting emotionally vulnerable individuals or causing stigmatization. To mitigate these risks, we recommend strict access controls, limiting system use to licensed mental health professionals. Clear documentation of the system’s capabilities and limitations is crucial to ensure it is used responsibly as a supportive aid.
Broader Societal Impact.
While our system aims to augment mental healthcare, we acknowledge that technology alone cannot solve systemic healthcare disparities. It should be considered one component within broader initiatives to expand mental health resources. We also note that this paper contains descriptions of depressive symptoms that may be distressing for some readers. We are committed to the ongoing ethical evaluation of our system and encourage community discussion on the responsible deployment of such technologies. Further discussion on the practical design considerations for deployment can be found in Appendix E.
Acknowledgement
This work was supported in part by the National Key Research and Development Program of China under grant 2022YFB3104702, in part by the China National Natural Science Foundation of China under grant 72442026, 72342032 and 62272262.
Appendix A Related Works
A.1 Depression Detection on Social Media
Compared to traditional medical diagnostic methods for depression conducted in hospitals, depression detection on social media has the advantages of lower concealment potential, wider coverage, and lower cost (Malhotra and Jindal, 2022). Early research works first extract text features, then apply traditional machine learning methods for classification. Common feature extraction methods in depression detection research include LIWC (Tausczik and Pennebaker, 2010), TF-IDF, LDA (Blei et al., 2003), etc.; classifiers include SVM (Tadesse et al., 2019), Logistic Regression (Coppersmith et al., 2014), Random Forest (Cacheda et al., 2019), among others.
In recent years, deep learning technologies have seen wide application in depression detection. Methods utilizing CNN (Cohan et al., 2018) and RNN (Ive et al., 2018) have been developed to capture complex patterns in text. Pre-trained Language Models (PLMs) (Owen et al., 2020; Ji et al., 2023) have further improved performance by leveraging knowledge learned from large corpora. Graph Convolutional Networks (GCNs) (Naseem et al., 2022) have also been employed to model user interactions and behaviors. Notably, AMM-Net (Sarkar et al., 2022) introduced an a multi-task learning model that combines word embeddings and topic features to effectively predict depression and anxiety on Reddit. Follow-up work extended this approach using Graph-based Message-passing Multi-task learning (GMMTL) (Sarkar et al., 2023). These efforts have significantly improved the accuracy of depression detection, advancing the field’s research.
Some works on depression detection focus solely on analyzing individual texts to determine the presence of depressive symptoms (Yang et al., 2022; Chiong et al., 2021), the type of symptoms displayed (Zhang et al., 2023a), and the level of depression (Zhang et al., 2023b). However, analyzing a user’s post history is more informative, as it is common for individuals not suffering from depression to also occasionally post texts that exhibit depressive symptoms. Recent approaches to depression detection based on a user’s post history Eichstaedt et al. (2018); Orabi et al. (2018); Lin et al. (2020); Chiong et al. (2021); Yang et al. (2022); Sarkar et al. (2022) lack comprehensive integration of medical knowledge, limiting their accuracy and explainability. While some studies have attempted to introduce medical knowledge by considering clinical symptoms (Yadav et al., 2020; Nguyen et al., 2022), they do so through relatively ineffective methods such as keyword matching or vector similarity matching, which may miss complex expressions of symptoms. Our research stands out as one of the initial efforts to more systematically apply medical insights, leading to improved accuracy. Additionally, we are pioneers in employing LLMs to craft systematic explanations of predictions grounded in clinical evidence, thereby increasing explainability.
A.2 Large Language Models
Large Language Models (LLMs), possessing rich common-sense knowledge and human-like reasoning abilities, have achieved tremendous success recently. LLMs are being applied to a wide variety of domains, such as mathematics and reasoning Xu et al. (2025); Hao et al. (2025b, a); HAO et al. (2024), text analysis Hao et al. (2024); Lan et al. (2024b, a), user behavior modeling Lan et al. (2025a, b); Feng et al. (2024); Du et al. (2024), user simulation Gao et al. (2023); Piao et al. (2025); Gao et al. (2024); Yan et al. (2025), code generation Jimenez et al. (2023); Hou et al. (2024); Ross et al. (2023), and embodied intelligence Brohan et al. (2023); Wang et al. (2025); Mon-Williams et al. (2025). Trained on vast amounts of human-generated text, LLMs have a strong potential to understand people, making them suitable for tasks like user behavior modeling and text comprehension. Consequently, they are also likely well-suited for tasks that require a deep understanding of human inner states, such as mental health analysis. Our work is among the first to explore this direction.
A.3 LLMs for Mental Health Analysis
Large language models (LLMs) have shown great potential in clinical applications due to its abundant prior knowledge and strong language generalization ability (Singhal et al., 2023). Recent works have introduced LLMs into mental health analysis. Arcan et al. (2024) presents a solid evaluation on the performance of Llama-2 (Touvron et al., 2023) and ChatGPT Ouyang et al. (2022) in mental health assessment tasks, unveiling prospects along with challenges of LLM-based methods. Yang et al. (2023a) further discover the applications of LLMs in both mental health detection tasks and reasoning tasks, which highlight LLM’s excellent interpretability. Yang et al. (2023b) propose the first open-source LLM series for mental health analysis based on Llama-2 and fine-tuning techniques, and greatly enhances the accuracy and explanation quality compared with general-purpose LLMs. However, the challenge of LLM-based methods being weaker in terms of prediction accuracy compared to embedding model-based methods, especially when targeting specific downstream tasks, remains unresolved (Arcan et al., 2024).
To our knowledge, our system is among the first to be specifically designed for depression detection in the context of LLM for mental health analysis. Furthermore, we have facilitated a collaboration between LLMs and embedding models, attaining both high accuracy and explainability, addressing deficiencies in previous designs.
Appendix B Datasets
B.1 SWDD Dataset
Acquiring well-labeled, large-scale datasets for depression detection is challenging due to the sensitivity and privacy concerns associated with mental illnesses. Some datasets utilize self-reported medical diagnoses. While this approach offers relatively high reliability, it leads to datasets comprising users who are already medically diagnosed and willing to discuss it online. However, the most valuable aspect of the depression detection task lies in assisting human experts to identify potentially undiagnosed users. Furthermore, the ratio of depressed to normal users in most previous datasets Losada and Crestani (2016); Zogan et al. (2021); Wu et al. (2023) is around 1:4 or even 1:1, deviating far from the actual prevalence of about 5% in the population.
To address these issues, we utilize the SWDD dataset Cai et al. (2023). This dataset carefully adheres to strict ethical guidelines. It includes users who self-disclosed a diagnosis and those identified by experts as highly likely to be depressed based on medical criteria, but without self-disclosure. We retain these non-self-disclosing users and, after further expert screening, keep 1000 depressed users and 19000 normal users to simulate a real-world application scenario with a 1:19 ratio. We retain only the posts posted within six months prior to each user’s last post. The dataset statistics are shown in Table 5. Training, validation, and test sets were divided in a 7:1:2 ratio, as recommended by Wu et al. (2023).
B.2 Twitter Mental Disorder Dataset
To further validate the generalizability of our approach, we also utilized the Twitter Mental Disorder Tweets and Musics dataset. This dataset contains tweets from users diagnosed with various mental disorders, including depression, alongside a control group. We utilized the subset relevant to depression detection for our experiments.
Table 5: Statistics of the SWDD dataset.
@@ B.2 Twitter Mental Disorder Dataset
Appendix C Details of Comparison Methods
Here, we give a more detailed discussion of our utilized comparison methods.
  • TF-IDF+XGBoost (Ramos et al., 2003; Chen and Guestrin, 2016; Wu et al., 2023): It uses TF-IDF weighted features of word and character n-grams for feature extraction, followed by XGBoost for classification.
  • HAN (Yang et al., 2016): It obtains tweet-level representations through bidirectional GRU networks, and encodes all representations into user presentation with an attention mechanism.
  • Mood2Content (Cai et al., 2023): This method utilizes knowledge distillation technology, transferring knowledge from an emotion classifier to depression detection.
  • AMM-Net (Sarkar et al., 2022): An attention-based multi-view and multi-task learning framework that integrates different feature perspectives and jointly optimizes for related mental health conditions (e.g., depression and anxiety).
  • FastText (Joulin et al., 2016), gte-small Li et al. (2023), BERT (Devlin et al., 2018), MentalRoBERTa (Ji et al., 2023): These models are used to obtain text embeddings. We compute the average embedding across all posts to generate a user-level representation, which is then fed into a tree classifier.
  • PHQ9 (Score) and PHQ9 (Vector) Nguyen et al. (2022): PHQ9 (Score) uses BERT to classify the 9 symptoms of PHQ-9, and then uses these symptom scores as features input to a CNN classifier. PHQ9 (Vector) uses the hidden layer vectors of the PHQ-9 symptom classifiers as features.
  • GPT-4o mini, MentalLLaMA (Yang et al., 2023b): For LLM-based approaches, we design prompts following Yang et al. (2023a). We input the user’s tweet history to the LLM, and then ask the LLM to classify.
Appendix D Hyperparameter Study (RQ4)
(a) Impact of $k$.
(b) Impact of $m$.
(c) Impact of $\alpha$.